iYuko (iyuko):

Introduction to Factoring! Come Learn with Me! Lesson 1

iYuko (iyuko):

Recall that when 2 or more numbers, variables, or algebraic expressions are multiplied, each is called a factor When you are asked to factor a number or a algebraic expression, you are being asked, "What, when multiplied, will give you that number or expression?" For example, you can factor 6 as 3*2 since 3*2=6. You can factor 15x^5 as 3x^2 * 5x^3 since 3x^2 * 5x^3 =15x^5. Factoring is simply the reverse of multiplying. 6 and 15x^5 are simple expressions to factor and can be factored in many ways. The factors of the polynomial \[2x ^{3}+x-12\] are not east to recognize In this tutorial we will be learning the techniques to find the factors of a polynomial. We will begin with simple factoring. When we factor, we are looking for the largest common factor. For example, in the polynomial 48x-16y, a common factor is 2. We could factor 48-16y as 2(24-8y). However, this is not complete. To factor 48-16y completely, we look for the greatest common factor of 48 and 16. 48-16y= 16(3x-y)

iYuko (iyuko):

Common factors of a polynomial 1. You can determine the numerical common factor by asking, "What is the largest integer that will divide into the coefficient of each term?" 2. You can determine the variable common factor by asking, "What variables are common to each term, and what is the largest exponent of those variables that is common?"

iYuko (iyuko):

Lets try an example: Factor. \[8x ^{3}y+16x ^{2} y ^{2}+24x ^{3} y ^{3}\] We see that 8 is the largest integer that will divide equally into the 3 numerical coefficients. We can factor x2 out of each term (not just x). For the other variable we factor y out of each term. \[8x ^{3} y+16x ^{2} y ^{2}+24x ^{3} y ^{3}=8x ^{2} y(x+2y+3yx ^{2})\] To check this you multiply the factors you obtain. The result should be the same as the original polynomial.

iYuko (iyuko):

Original post is at: http://openstudy.com/users/iyuko#/updates/552f38fae4b01573da53b3cf

TheSmartOne (thesmartone):

Now I know how to factor! :D Thanks :)

OpenStudy (anonymous):

Wow great job! @iYuko

OpenStudy (anonymous):

nice job

iYuko (iyuko):

Thanks @TheSmartOne @Tom_Boy_Rebel @peak

OpenStudy (alexandervonhumboldt2):

nice job :)

OpenStudy (anonymous):

help me Point A is located at negative 7 over 10 and point B is located at negative 2 over 10. What is the distance between points A and B?

iYuko (iyuko):

Thank you. I made this a long time ago, I thought I should re post it.

iYuko (iyuko):

I'm bout to re post another.

iYuko (iyuko):

You can find these tutorials at http://openstudy.com/users/iyuko#/updates/555e5f3ee4b0dbf81b467f4a

OpenStudy (anonymous):

thanks

TheSmartOne (thesmartone):

Are you just re-posting to get more medals? :P

TheSmartOne (thesmartone):

Uhh wait, of course!

iYuko (iyuko):

Well, duh! Plus people need help on it ;)

iYuko (iyuko):

It's a great tutorial. People need to know how to do it.